Two point charges are located on the y-axis of a  coordinate system: q1 = −3.0 nC is at y = −2.0 cm, and q2 = 1.0 nC is at y = −4.0 cm. What is the total electric force exerted by q1 and q2 on a charge q3 = 5.0 nC at y = 0?


1
Expert's answer
2021-02-08T18:40:05-0500

As per the given question,

"q_1=-3.0nC=3\times 10^{-9}C"

position co-ordinate "y_1=-2.0cm=-0.02m"

"q_2=1\times 10^{-9}C"

"y_2=-4.0cm = -0.04m"

"q_3=5\times 10^{-9}C"

"y_3=0"

So, net force on the charge "q_3"


"F_{net}=\frac{q_1q_3}{4\pi (y_3-y_1)^2}-\frac{q_2q_3}{4\pi (y_3-y_2)^2}"


Now, substituting the values,


"F_{net}=\frac{3\times 10^{-9}\times5\times10^{-9}\times 9\times 10^9}{(0.02-0)^2}+\frac{1\times 10^{-9}\times5\times10^{-9}\times 9\times 10^9}{(0.04-0)^2}"


"=\frac{135\times 10^{-9}}{4\times 10^{-4}}+\frac{45\times 10^{-9}}{16\times 10^{-4}}"


"=\frac{(540+45)\times 10^{-5}}{16}N"


"=\frac{595\times 10^{-5}}{16}N"


"=37.18\times 10^{-5}N"